# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### 12 Jan 2024 v8

@@ 5.2. FM adapted to DMs through representation-level knowledge transfer
PrivateLoRA (Wang et al., 2023a) Forward Repr, Backward gradient Forward Repr, Backward gradient
@@
In a two-party federated split learning setting involving a client and a server, the client typically owns a private dataset and a small portion of an FM, while the server owns the larger portion. The objective of FM-based split learning is to adapt the knowledge of FM to the client’s specific domain by fine-tuning the distributed FM using the client’s local data. The FM is often split into an FM backbone and multiple other smaller parts. The FM backbone is the most computationally expensive part and is typically deployed on a server with sufficient computing resources, while other parts can be deployed on either the client or server. The training procedure of distributed FM follows the conventional split learning procedure using stochastic gradient descent, but each portion of FM is trained by a different party (i.e., client and server). Figure 8 illustrates three two-party federated split learning settings studied in the literature.
@@
PrivateLoRA (Wang et al., 2023a) explored the third setting of federated split learning, as illustrated in Figure 8(c). Inspired by LoRA, PrivateLoRA represents the update $\Delta W$ on each linear projection of query, key, or value in self-attention with three sequential low-rank matrices $\Delta W=A\times M\times B$. Consequently, there are $\{A\}_{i=1}^{K}$, $\{B\}_{i=1}^{K}$, and $\{M\}_{i=1}^{K}$ for $K$ target linear projections. PrivateLoRA deploys $\{A\}_{i=1}^{K}$ and $\{B\}_{i=1}^{K}$ on the server, whereas $\{M\}_{i=1}^{K}$ on the client in addition to deploying the FM backbone on the server and FM head and bottom models on the client. The $k$th FM middle model in Figure 8(c) corresponds to $M_{k}$. During training, $\{A\}_{i=1}^{K}$$\{B\}_{i=1}^{K}$, and the FM backbone are frozen, while the $\{M\}_{i=1}^{K}$, FM head, and FM bottom models are optimized. The benefits of PrivateLoRA lie in enhancing the performance of the distributed FM on the client’s local task while significantly diminishing communication overheads.
@@ 5.3. FM adapted to DMs through model-level knowledge transfer
For homo-model transfer without adaptation, JFT (Shi et al., 2022b) proposed a selective differential privacy approach to improve the utility of DP-tuned FMs (i.e., Roberta-base and GPT2-small), which can be released to downstream clients. More specifically, JFT first fine-tunes an FM with the redacted version of a private dataset and then fine-tunes the FM again with the original private data using an efficient DP-SGD (Li et al., 2022c). DP-BiTFiT (Bu et al., 2022) aims to improve the efficiency of fine-tining LLMs with DP-SGD by training only bias terms in an FM.  Youssef et al. (2023) proposed a privacy-preserving approach to share an FM (i.e., $\text{BERT}_{\text{base}}$) pre-trained on a private dataset. More specifically, they proposed to protect the privacy of the private data by sharing only parts of the FM’s pretrained parameters with the rest parameters randomly initialized and releasing a proxy of the private data for downstream clients to fine-tune the FM for complementary knowledge transfer.
@@
For transferring hetero-models adaptive to downstream tasks,  Wang et al. (2023b) proposed an approach that leverages public data and pre-trained FM (i.e., LaMDA (Cohen et al., 2022)) to enhance downstream cross-device federated learning. More specifically, they first leverage their proposed distribution matching algorithm to sample public data closely resembling the distribution of FL clients’ private data, and then they use the sampled data to distill a pre-trained FM (i.e., LaMDA) into an on-device language model, which serves as an initialization for FL clients. GPT-FL (Zhang et al., 2023b) leverages FMs (i.e., Stable Diffusion (Rombach et al., 2022), SpeechT5 (Ao et al., 2022), and AudioLDM (Liu et al., 2023a)) to generate diversified synthetic data to train a downstream model, which is then distributed to FL clients for model initialization.
@@ 5.4. Privacy and Backdoor Attacks
Privacy attacks focus on inferring private data from knowledge at the representation level. More specifically, TextObfuscator (Zhou et al., 2023), RAPT (Li et al., 2023c) and Xu et al. (2023c) investigated model inversion attacks in federated split learning scenarios, in which the server aims to reconstruct the client’s private data based on the representations output from the client’s local model. These studies proposed representation perturbation, $d_{\chi}$-privacy, and weight permutation equivalence, respectively, to thwart model inversion attacks. The rationale behind these defenses is to reduce the correlation between private data and the representations shared with the server. As a result, the chance for the adversary to recover private data can be reduced.
@@
6. FTL-FM: Domain Knowledge for Augmenting Foundation Models
@@ 6.3. Privacy and Backdoor Attacks
7. FTL-FM: Co-optimize Foundation Models and Domain Models
@@ 7.1. Co-optimize FMs and DMs through model-level knowledge transfer
    The server serves as an aggregation function while clients’ domain models are initialized with or assisted by FMs. The objective of this scenario is to fine-tune clients’ DMs through federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
@@
For the first scenario, FATE-LLM (Fan et al., 2023) proposed a FedCoLLM (Federated Co-tuning LLM) framework (Fan et al., 2023) to co-optimize the server’s FM and clients’ DMs simultaneously through mutual distillation. Figure 12(a) illustrates the workflow of FedCoLLM. In FedCoLLM, clients’ DMs are initialized with an off-the-shelf FM (i.e., LLaMA-7B (Touvron et al., 2023)). In each communication round, clients fine-tune their DMs using local data and then send them to the server with secure aggregation (Bonawitz et al., 2016). The server, in turn, distills the knowledge mutually between its hosted FM (i.e., LLaMa-65B) and the global DM aggregated from clients’ local DMs to co-optimize both models. Subsequently, the server dispatches the global DM to all clients for further training.  Deng et al. (2023) proposed CrossLM that optimizes the server’s FM and clients’ DMs through synthetic data that are generated by the server’s FM and validated by clients’ DMs, as illustrated in Figure 12(b). Specifically, in each communication round, clients train their DMs using local data and then send them to the server asynchronously. The server, in turn, generates synthetic data and leverages received DMs to validate the quality of synthetic data. Subsequently, the server utilizes validation feedback from DMs and the validated synthetic data to fine-tune the server’s FM. Next, the server uses the validated synthetic data to fine-tune each client’s DM and dispatches the fine-tuned DMs to corresponding clients for further training. This concerted effort is aimed at co-optimizing both the FM and DMs. FedCoLLM and CrossLM adhere to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedCoLLM has $\eta>0$ whereas CrossLM has $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and a client $k$’s $\mathcal{W}^{C_{k}}$ is formulated in Eq.(12).
@@
As for now, few works following Eq.(5) explicitly consider protecting data privacy (i.e., $\eta>0$). FedPETuning (Zhang et al., 2023d), PEU+LoRA (Xu et al., 2023a), DP-LoRA (Liu et al., 2023d), and FedPrompt (Zhao et al., 2023b) are representative ones. FedPETuning leverages four PEFT techniques, namely adapter tuning, prefix tuning, LoRA (Hu et al., 2022), and BitFit (Ben Zaken et al., 2022), to fine-tune clients’ DMs (i.e., $\text{Roberta}_{\text{base}}$) and demonstrated that federated learning combined with LoRA achieved the best privacy-preserving results. DP-LoRA and PEU+LoRA apply LoRA with DP to fine-tune clients’ DMs, aiming to achieve better privacy-utility-resource trade-offs than baselines. PEU+LoRA additionally uses Partial Embedding Updates (PEU) to reduce the noise required to protect privacy. Zhao et al. (2023b) proposed FedPrompt that leverages federated learning combined with soft prompt tuning to fine-tune clients’ DMs (e.g., T5 (Raffel et al., 2020)), aiming to reduce communication cost while maintaining model accuracy. FedPrompt adopts local DP to protect the privacy of clients’ data, demonstrating there is a trade-off between privacy and model accuracy.
@@ 8. Foundation Model Inference
When the FM provider is the adversary, its goal is to infer private information of the user by investigating the prompts sent by the user with the assistance of the FM. The FM provider can infer privacy information from the prompt of a single query (illustrated in Figure 14(a)) (Tong et al., 2023), or it can steer a seemingly benign conversation with the user to lure the user into providing prompts that can be exploited to learn private and sensitive information 14(b)) (Tong et al., 2023).
In literature, most of the privacy-preserving methods that aim to protect the user’s data privacy during inference focus on designing efficient Secure Multi-Party Computation (SMPC) methods (Li et al., 2023d; Dong et al., 2023; Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023). Given that SMPC comes at a significant cost in terms of computational resources and communication, these methods primarily focus on optimizing the efficiency of FM architectures, predominantly Transformer-based neural networks, or SMPC protocols. We will comprehensively review these SMPC-based methods in Section 10. In a recent study, Tong et al. (2023) proposed InferDPT, which leverages DP-guaranteed text perturbation on the user’s prompts to prevent the adversary FM provider from inferring the user’s private data. Specifically, in the InferDPT framework, the user first randomly perturbs tokens of its prompt with semantically similar alternatives. Subsequently, the user sends the perturbed prompt to an LLM (i.e., GPT-4), which in turn generates a corresponding response and sends the response back to the user. Finally, the user employs a local pre-trained language model, such as Vicuna-7B, to generate the final response based on the original prompt and the response received from the LLM.  Staab et al. (2023) investigated leveraging text anonymization to protect the user’s data privacy, and it empirically demonstrated that anonymization is insufficient for protecting the user’s data privacy.
@@
When the user is the adversary, his/her goal is to extract the private information of some targeted people from the FM (illustrated in Figure 15(a)) (Carlini et al., 2021) or from the private data sources used to augment the FM (illustrated in Figure 15(b)) through privacy-invasive prompts (Panda et al., 2023).
For defending against privacy-invasive prompts, Panda et al. (2023) proposed a DP-ICL approach that allows the LLM to generate differentially private responses through a noisy consensus among an ensemble of responses. Staab et al. (2023) suggested that model alignment can be used to thwart privacy-invasive prompting. Nonetheless, they also demonstrated that current LLMs are not aligned against privacy-invasive prompts because much of the current alignment research focuses primarily on preventing the LLM from generating harmful and offensive content. Thus, leveraging alignment to safeguard the FM provider’s privacy during inference presents a promising avenue for further research.

### 07 Feb 2024 v10

@@ 5.2. FM adapted to DMs through representation-level knowledge transfer
PrivateLoRA (Wang et al., 2023b) Forward Repr, Backward gradient Forward Repr, Backward gradient
@@
In a two-party federated split learning setting involving a client and a server, the client typically owns a private dataset and a small portion of an FM, while the server owns the rest larger portion. The objective of FM-based split learning is to adapt the knowledge of FM to the client’s specific domain by fine-tuning the distributed FM using the client’s local data. The FM is often split into an FM backbone and multiple other smaller parts. The FM backbone is the most computationally expensive part and is typically deployed on a server with sufficient computing resources, while other parts can be deployed on either the client or server. The training procedure of distributed FM follows the conventional split learning procedure using stochastic gradient descent, but each portion of FM is trained by a different party (i.e., client and server). Figure 8 illustrates three two-party federated split learning settings studied in the literature.
@@
SAP (Shen et al., 2023) takes a similar setting as TextObfuscator, but it adopts a different training strategy. During training, the client model is frozen, while the server model is trained using LoRA. To protect its data privacy, the client applies $d_{\chi}$-privacy (Chatzikokolakis et al., 2013) to obfuscate representation vectors before transmitting them to the server. Furthermore, SAP studied the impact of the FM’s split position on the privacy-preserving capacity and the model performance of the SAP framework. It empirically demonstrated that as the number of layers in the bottom model increases, the utility diminishes while the privacy-preserving capability strengthens, highlighting the existence of a trade-off between utility and privacy when partitioning the FM. SAP is a variant of the typical federated split learning we formulated in Eq.(8) as it conducts the FM adaptation through fine-tuning the model on the server side, as opposed to the client model.
PrivateLoRA (Wang et al., 2023b) explored the third setting of federated split learning, as illustrated in Figure 8(c). Inspired by LoRA, PrivateLoRA represents the update $\Delta W$ on each linear projection of query, key, or value in self-attention with three sequential low-rank matrices $\Delta W=A\times M\times B$. Consequently, there are $\{A\}_{i=1}^{K}$, $\{B\}_{i=1}^{K}$, and $\{M\}_{i=1}^{K}$ for $K$ target linear projections. PrivateLoRA deploys $\{A\}_{i=1}^{K}$ and $\{B\}_{i=1}^{K}$ on the server, whereas $\{M\}_{i=1}^{K}$ on the client in addition to deploying the FM backbone on the server and FM head and bottom models on the client. The $k$th FM middle model in Figure 8(c) corresponds to $M_{k}$. During training, $\{A\}_{i=1}^{K}$$\{B\}_{i=1}^{K}$, and the FM backbone are frozen, while the $\{M\}_{i=1}^{K}$, FM head, and FM bottom models are optimized. The benefits of PrivateLoRA lie in enhancing the performance of the distributed FM on the client’s local task while significantly diminishing communication overheads.
@@ 5.3. FM adapted to DMs through model-level knowledge transfer
For homo-model transfer without adaptation, JFT (Shi et al., 2022b) proposed a selective differential privacy approach to improve the utility of DP-tuned FMs (i.e., Roberta-base and GPT2-small), which can be released to downstream clients. More specifically, JFT first fine-tunes an FM with the redacted version of a private dataset and then fine-tunes the FM again with the original private data using an efficient DP-SGD (Li et al., 2022d). DP-BiTFiT (Bu et al., 2022) aims to improve the efficiency of fine-tining LLMs with DP-SGD by training only bias terms in an FM.  Youssef et al. (2023) proposed a privacy-preserving approach to share an FM (i.e., $\text{BERT}_{\text{base}}$) pre-trained on a private dataset. More specifically, they proposed to protect the privacy of the private data by sharing only parts of the FM’s pretrained parameters with the rest parameters randomly initialized and releasing a proxy of the private data for downstream clients to fine-tune the FM for complementary knowledge transfer.
@@
For transferring hetero-models adaptive to downstream tasks,  Wang et al. (2023c) proposed an approach that leverages public data and pre-trained FM (i.e., LaMDA (Cohen et al., 2022)) to enhance downstream cross-device federated learning. More specifically, they first leverage their proposed distribution matching algorithm to sample public data closely resembling the distribution of FL clients’ private data, and then they use the sampled data to distill a pre-trained FM (i.e., LaMDA) into an on-device language model, which serves as an initialization for FL clients. GPT-FL (Zhang et al., 2023b) leverages FMs (i.e., Stable Diffusion (Rombach et al., 2022), SpeechT5 (Ao et al., 2022), and AudioLDM (Liu et al., 2023a)) to generate diversified synthetic data to train a downstream model, which is then distributed to FL clients for model initialization.
@@ 5.4. Privacy and Backdoor Attacks
Privacy attacks focus on inferring private data from knowledge at the representation level. More specifically, TextObfuscator (Zhou et al., 2023a), RAPT (Li et al., 2023c) and Xu et al. (2023c) investigated model inversion attacks in federated split learning scenarios, in which the server aims to reconstruct the client’s private data based on the representations output from the client’s local model. The three studies proposed representation perturbation, $d_{\chi}$-privacy, and weight permutation equivalence, respectively, to thwart model inversion attacks. The rationale behind these defenses is to reduce the correlation between private data and the representations shared with the server. As a result, the chance for the adversary to recover private data can be reduced.
@@
6. Domain Knowledge for Augmenting Foundation Models
@@ 6.3. Privacy and Backdoor Attacks
7. Co-optimize Foundation Models and Domain Models
@@ 7.1. Co-optimize FMs and DMs through model-level knowledge transfer
    The server serves as an aggregation function while clients’ DMs are initialized with or assisted by FMs. The objective of this scenario is to fine-tune to train clients’ DMs through federated learning. Each client ends up with a DM enhanced with the domain knowledge of all clients.
For the first scenario, FATE-LLM (Fan et al., 2023) proposed a FedCoLLM (Federated Co-tuning LLM) framework (Fan et al., 2023) to co-optimize the server’s FM and clients’ DMs simultaneously through mutual distillation. Figure 12(a) illustrates the workflow of FedCoLLM. In FedCoLLM, clients’ DMs are initialized with an off-the-shelf FM (i.e., LLaMA-7B (Touvron et al., 2023)). In each communication round, clients fine-tune their DMs using local data and then send them to the server with secure aggregation (Bonawitz et al., 2016). The server, in turn, distills the knowledge mutually between its hosted FM (i.e., LLaMa-65B) and the global DM aggregated from clients’ local DMs to co-optimize both models. Subsequently, the server dispatches the global DM to all clients for further training.
Deng et al. (2023) proposed CrossLM that optimizes the server’s FM and clients’ DMs through synthetic data that are generated by the server’s FM and validated by clients’ DMs, as illustrated in Figure 12(b). Specifically, in each communication round, clients train their DMs using local data and then send them to the server asynchronously. The server, in turn, generates synthetic data and leverages received DMs to validate the quality of synthetic data. Subsequently, the server utilizes validation feedback from DMs and the validated synthetic data to fine-tune the server’s FM. Next, the server uses the validated synthetic data to fine-tune each client’s DM and dispatches the fine-tuned DMs to corresponding clients for further training. This concerted effort is aimed at co-optimizing both the FM and DMs. FedCoLLM and CrossLM adhere to the objective formulated in Eq.(4) with $K>1$, $\sum^{K}_{k=1}p^{C_{k}}=1$, and $p^{S}=0$. FedCoLLM has $\eta>0$ whereas CrossLM has $\eta=0$. The specific form of the loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};d)$ for co-optimizing the server’s $\mathcal{W}^{S}_{\text{FM}}$ and a client $k$’s $\mathcal{W}^{C_{k}}$ is formulated in Eq.(3.3).
@@
As for now, few works following Eq.(5) explicitly consider protecting data privacy (i.e., $\eta>0$). FedPETuning (Zhang et al., 2023d), PEU+LoRA (Xu et al., 2023a), DP-LoRA (Liu et al., 2023d), FedPrompt (Zhao et al., 2023b), and the work of Huang et al. (2024) are representative ones. FedPETuning leverages four PEFT techniques, namely adapter tuning, prefix tuning, LoRA (Hu et al., 2022), and BitFit (Ben Zaken et al., 2022), to fine-tune clients’ DMs (i.e., $\text{Roberta}_{\text{base}}$) and demonstrated that federated learning combined with LoRA achieved the best privacy-preserving results. DP-LoRA and PEU+LoRA apply LoRA with DP to fine-tune clients’ DMs, aiming to achieve better privacy-utility-resource trade-offs than baselines. PEU+LoRA additionally uses Partial Embedding Updates (PEU) to reduce the noise required to protect privacy. Zhao et al. (2023b) proposed FedPrompt that leverages federated learning combined with soft prompt tuning to fine-tune clients’ DMs (e.g., T5 (Raffel et al., 2020)), aiming to reduce communication cost while maintaining model accuracy. FedPrompt adopts local DP to protect the privacy of clients’ data, demonstrating there is a trade-off between privacy and model accuracy. Huang et al. (2024) proposed a method that involves fine-tuning clients’ DMs using LoRA and P-Tuning v2. A key consideration in their approach is ensuring the privacy of both clients’ data and models. To address this concern, they implemented the fine-tuning process within a Trusted Execution Environment (TEE). Additionally, they employed One Time Pad (OTP) to secure all communicated messages.
@@ 8. Foundation Model Inference
Table 8. Summary of adversaries, privacy attacks, protection methods of foundation model inference. EIA: Embedding Inversion Attack; AIA: Attribute Inference Attack; LIA: LLM-assisted Inference Attack; PIP: Privacy-Invasive Prompt; SMPC: Secure Multi-Party Computation; DP: Differential Privacy.
Adversary Attacking Target Attack Method Protection Method
FM Provider Private information conveyed in the prompts of the user EIA (Qu et al., 2021; Song and Raghunathan, 2020; Morris et al., 2023; Kugler et al., 2021) AIA (Plant et al., 2021; Li et al., 2022c; Song and Shmatikov, 2019; Hayet et al., 2022; Yue et al., 2021) LIA (Tong et al., 2023; Staab et al., 2024) SMPC (Li et al., 2023d; Dong et al., 2023; Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023) DP-based perturbation (Tong et al., 2023) DP-based ensemble (Hong et al., 2023) Anonymization (Staab et al., 2024)
FM User Private data used to train or augment the FM PIP (Panda et al., 2023; Staab et al., 2024; Carlini et al., 2021) DP-based ensemble (Panda et al., 2023)
Filtering (Perez et al., 2022)
Alignment (Staab et al., 2024)
@@
When FM providers are adversaries, they may attempt to infer the private information of the user by investigating the prompts sent by the user with the assistance of the FM. The adversarial FM provider can infer privacy information from the prompt of a single query (illustrated in Figure 14(a)) (Tong et al., 2023), or it can steer a seemingly benign conversation with the user to lure the user into sending out prompts that can be exploited to learn private and sensitive information (illustrated in Figure 14(b)) (Staab et al., 2024). Various attacking methods have been proposed in these scenarios. For example, FM providers can carry out embedding inversion attacks to reconstruct the original input text based on the provided embedding (Qu et al., 2021; Song and Raghunathan, 2020; Morris et al., 2023; Kugler et al., 2021). They can also employ attribute inference attacks to deduce specific attributes such as race, gender, and age from the given embedding or text (Plant et al., 2021; Li et al., 2022c; Song and Shmatikov, 2019; Hayet et al., 2022; Yue et al., 2021). The FM provider can also leverage its LLM to recover the original input from its perturbed version (Tong et al., 2023; Staab et al., 2024).
In literature, most of the privacy-preserving methods that aim to protect the user’s data privacy during inference focus on designing efficient Secure Multi-Party Computation (SMPC) methods (Li et al., 2023d; Dong et al., 2023; Liu and Liu, 2023; Zheng et al., 2023; Hou et al., 2023a; Hao et al., 2022; Ding et al., 2023). Given that SMPC comes at a high cost in terms of computational resources and communication, these methods primarily focus on optimizing the efficiency of FM architectures, predominantly Transformer-based neural networks, or SMPC protocols. We will comprehensively review these SMPC-based methods in Section 10. In a recent study, Tong et al. (2023) proposed InferDPT, which leverages DP-guaranteed text perturbation on the user’s prompts to prevent the adversary FM provider from inferring the user’s private data. Specifically, in the InferDPT framework, the user first randomly perturbs tokens of its prompt with semantically similar alternatives. Subsequently, the user sends the perturbed prompt to an LLM (i.e., GPT-4), which in turn generates a corresponding response and sends the response back to the user. Finally, the user employs a local pre-trained language model, such as Vicuna-7B, to generate the final response based on the original prompt and the response received from the LLM. DP-OPT (Hong et al., 2023) adopts Deep Language Network (DLN (Sordoni et al., 2023)) guided by a local FM to automatically optimize prompts. During the optimization, DP-OPT leverages a differentially private ensemble method to generate prompts to protect prompts’ privacy.  Staab et al. (2024) investigated leveraging text anonymization to protect the user’s data privacy, and it empirically demonstrated that anonymization is insufficient for protecting the user’s data privacy.
When users are adversaries, their goal is to extract the private information of some targeted people from the FM (illustrated in Figure 15(a)) (Carlini et al., 2021; Staab et al., 2024) or from the private data sources used to augment the FM (illustrated in Figure 15(b)) (Panda et al., 2023) through designing privacy-invasive prompts.
For defending against privacy-invasive prompts, Panda et al. (2023) proposed a DP-ICL approach that allows the LLM to generate differentially private responses through a noisy consensus among an ensemble of responses.  Perez et al. (2022) proposed to remove sensitive text from the response of the FM before it is sent to users, aiming to prevent the FM’s response from leaking private information. Staab et al. (2024) suggested that model alignment can be used to thwart privacy-invasive prompting. Nonetheless, they also demonstrated that current LLMs are not aligned against privacy-invasive prompts because much of the current alignment research focuses primarily on preventing the LLM from generating harmful and offensive content. Thus, leveraging alignment to protect the FM provider’s privacy during inference presents a promising avenue for further research.